# A Large-Scale Benchmark of Cross-Modal Learning for Histology and Gene Expression in Spatial Transcriptomics

### v1

@@ 1 Introduction
Recent years have witnessed remarkable advances in representation learning for biomedicine, enabling novel strategies to integrate multimodal objectives [30, 12, 46]. Advancements include learning joint representations of whole-slide pathology images and clinical reports [29, 45], integration of pathology image data with bulk gene expression profiles [22], and combining x-ray scans with clinical notes [18]. These successes have been accelerated by the availability of large-scale, carefully curated datasets and robust training paradigms that leverage cross-modal information to yield richer, more generalizable features across domains.
@@
Unimodal foundation models have pushed the boundaries for representation learning in pathology [6, 48]. Several instances of large-scale foundation models are available and have demonstrated strong performance across a range of challenging benchmarks [59, 33, 25, 32, 4]. While unimodal foundation models in pathology have shown promise [6, 51, 57, 41, 29], fully harnessing spatial transcriptomics demands new multimodal strategies capable of learning from both tissue images and gene expression measurements simultaneously. Such approaches have the potential to highlight principles of tissue architecture that might remain obscured in single-modality analyses, as well as enhance model performance in unimodal predictive tasks.
To elucidate this challenge, we curated a large-scale dataset of over $\sim 720k$ patch-level image–gene pairs. These data span multiple organs and originate from 6 different 10x Xenium in-situ panels, collected across 54 patients and 57 individual samples. Filtering out replicates with different gene panels and single gene panel datasets, we use $\sim 620k$ image-gene pairs for the current benchmark. Through systematic examination of image and gene expression encoder combinations across a comprehensive set of configurations, we present an extensive benchmark of modality-specific encoder performance in contrastive pretraining settings. Our results demonstrate that dataset-specific pretraining of gene expression encoders has the greatest impact on cross-modality retrieval performance. Additionally, we emphasize the critical importance of addressing batch effects in these encoders. For the gene expression modality specifically, the ability to directly model batch effects provides significant performance improvements in cross-modality retrieval tasks.
@@ 3.3.1 Losses
Specifically, for a minibatch of images and gene expression pairs =$\mathcal{B}=\left\{\left(I_{1},G_{1}\right),\left(I_{2},G_{2}\right),\ldots\right\}$, the CLIP learning objective for the image-gene expression pair v2g is:
@@ 3.3.2 Metrics
[…]
@@ 3.4.2 HEST Benchmark
Fine-tuned image encoders are evaluated on the HEST benchmark [21] by regressing patch-level gene expression from image patches. The benchmark comprises nine tasks that span various tissue types and cancer subtypes from both 10x xenium and visium datasets. Each task predicts the expression of the top 50 most variable genes from 224$\times$224 pixel histology patches 112 $\mu$m $\times$ 112 $\mu$m at 20x magnification. The features extracted at the patch level are then passed to a ridge regression model for gene expression prediction.

### v2

@@ 1 Introduction
Recent advances in biomedical representation learning have enabled joint modeling of pathology images with reports [29, 46], bulk gene expression [22], and x-ray scans with clinical notes [18], largely driven by large-scale curated datasets and robust training paradigms [12, 30, 47].
@@
Pathology foundation models have achieved strong performance across benchmarks [6, 49, 60, 34, 25, 32, 4]. To fully harness the information provided by spatial transcriptomics, we require novel multimodal strategies capable of learning of integrating tissue images with gene expression.
To elucidate this challenge, we curated a large-scale dataset of over $\sim 620k$ patch-level image–gene pairs across multiple organs originating from 6 different 10x Xenium in-situ panels, collected across 54 patients and 57 individual samples. This dataset enable us to perform a large-scale benchmark of image and gene encoders using contrastive pretraining. Our results demonstrate that dataset-specific pretraining of gene expression encoders has the greatest impact on cross-modality retrieval performance. Additionally, we emphasize the critical importance of addressing batch effects in these encoders. For the gene expression modality specifically, the ability to directly model batch effects provides significant performance improvements in cross-modality retrieval tasks.
@@ 3.3.1 Losses
Specifically, for a minibatch of images and gene expression pairs $\mathcal{B}=\left\{\left(I_{1},G_{1}\right),\left(I_{2},G_{2}\right),\ldots\right\}$, the CLIP learning objective for the image-gene expression pair v2g is:
@@ 3.3.2 Metrics
We evaluate the representation quality after image-transcriptomics alignment with Recall@$k$ for $k=\{1,5,10\}$, which measures the proportion of relevant items retrieved within the top $k$ retrieved samples. We report both image-to-gene (I2G) retrieval, where an image query retrieves its corresponding gene-expression profile, and gene-to-image (G2I) retrieval, where a gene-expression query retrieves the matching histological image.
@@ 3.4.2 HEST Benchmark
Fine-tuned image encoders are evaluated on the HEST benchmark [21] by regressing patch-level gene expression from image patches. The benchmark comprises nine tasks that span various tissue types and cancer subtypes from both 10x xenium and visium datasets. Each task predicts the expression of the top 50 most variable genes from $\times224$ pixel histology patches $\mum ×\times 112 μ\mum$ at 20x magnification. The features extracted at the patch level are then passed to a ridge regression model for gene expression prediction.